OpenStudy (anonymous):

A group of friends were voting for their favorite song: two-fifths chose song A, one-sixth of the remaining friends chose song B, and the rest chose song C. If there were thirty friends in all, how many chose song C? ten twelve fifteen eighteen

OpenStudy (anonymous):

12 / 30 chose song A, 5/30 chose song B, and that leaves 13/30 to choose song C.

OpenStudy (anonymous):

scratch that im fixing it

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

15 chose song C. A= (times by 6)12/30 . Remove those 12 from he group because it says the REMAINING chose song B. total = 18 (atm) Song B= 3/18 (multiply 1/6 by 3) Song C= 15 people chose songs A + B so now theyre 15 people left who chose song C

OpenStudy (anonymous):

THANK YOU :) sorry for the caps